OpenStudy (anonymous):

A line with a slope of 1/4 passes through (-1,-5). Write the equation of the line in general form.

OpenStudy (anonymous):

x - 4y = 19 y = 1/4x - 19/4 x + 4y = 6 y + 5 = 1/4(x + 1)

OpenStudy (anonymous):

PLEASE HELP!!